# Robustly optimal dynamics for active matter reservoir computing

### v1

@@ Abstract
We study the information processing abilities of active matter in the reservoir computing (RC) paradigm, using a model that is externally driven to infer the future state of a chaotic signal. The simulated system closely follows a previously reported model. We uncover an exceptional dynamical regime of agent dynamics that has been overlooked heretofore. It appears robustly optimal across varying physical parameters and inference tasks, thus providing valuable insights into computation and inference with physical systems more generally. The ability to form effective mechanisms for information processing are primarily determined by the system’s own intrinsic relaxation abilities. These are identifiable when probing the system without a specific inference goal and manifest when testing minimalistic single-particle reservoirs. The regime that achieves optimal computation is situated just below the critical damping threshold, involving a microscopic dynamical relaxation with multiple stages. The optimal system is adaptable under chaotic external driving, due to a diversity in response mechanisms that emerge like rapid alternations between quasi-stationary and highly nonlinear dynamical states. Both coherent and incoherent dynamics contribute to their operation, partly at dissimilar scales of space and delay time. Correlations on agent dynamics can indicate the best-performing regimes and onsets of tight relationships between the responding system and the fluctuating driver. As this model of computation is interpretable in physical terms, it facilitates re-framing inquiries regarding learning and unconventional computing with a fresh rationale for many-body physics out of equilibrium.
@@ I Introduction
The reservoir computing (RC) paradigm is ideal for exploring these basic questions, as it exploits directly dynamical systems for information processing and inference (in real-time) [25, 26, 4]. The power of RC for scientific inquiry into physical computation is its generic and straightforward setup: Information of a low-dimensional input (e.g. a time series) is injected into the reservoir computing substrate, which is a high-dimensional nonlinear dynamical system that enters different dynamical states depending on the input signal and its history [27]. In the RC paradigm, the reservoir substrate is optimized based on heuristic global properties (such as the spectral radius in Echo State Networks). In the canonical case of using a digital neural network, substrates are simply initialized in a random way. The readout layer [28] typically consists of a linear model trained using Ridge regression. This separation between training the nonlinear substrate and a simple readout layer constitutes the original discovery of RC as a method [25, 26, 4], avoiding costly backpropagation through time [28].
RC renders state-of-the-art performance for time-series forecasting [29, 30, 31, 32], in particular, chaotic time series prediction [30]. Despite the plethora of other machine learning techniques available for time-series forecasting (nonlinear vector autoregression (“next-generation RC”) [33], NBEATS [34], NHiTS [35], transformer models [36] or LSTMs [37]), the highly generic method of RC using e.g. Echo State Networks has often remained an architecture of choice. It requires far less training data [31, 32] and allows for predictions several Lyapunov times ahead in a closed-loop autoregressive fashion [31]. It is considered “model-free” in the sense that it can estimate properties of the input without the need for an explicit model thereof [38]. It can be implemented using generic substrates, e.g., enabling highly competitive photonics-based machine learning [39, 40, 41].
@@
However, many questions remain unresolved: For example, what is the optimal mix of nonlinearity and linearity in responses? [46] Do memory capacity and nonlinearity present trade-offs? [47, 46] Is computation most effective near a critical point or at the “edge of chaos”? [48, 49, 50, 51, 52]; How can input information be effectively injected into the system for more efficient RC? [33] What about the readout? [53] And how can interpretable frameworks be developed to integrate theory with practice and enhance the prevalence of RC? [54, 33].
@@
Yet, for reservoir computation in materio, deeper physical concepts lie hidden beneath the surface. For example, how do nonequilibrium phase diagrams and dissipative as well as correlative behavior relate to computational properties? Understanding this fundamental connection would be valuable for recognizing the true feasibility of general neuromorphic computing systems [19, 55, 12], and connect different realms of matter-based and digital computing, both conceptually and practically [56]. Artificial neural networks are usually employed as the model of RC, to develop mathematical theory; they are not particularly interpretable in these terms, essentially functioning as black-boxes [57]. Thus, a different paradigm for thinking about computation and learning is required when dealing with physical substrates [1, 13].
@@
RC has been demonstrated for soft matter, though extremely rarely, and developments remain in an early stage: Reported were cases of gel materials, soft macroscopic bodies, and biological cell cultures [58, 59, 60]. In one exotic example, water waves in a bucket were used as a proof-of-principle [61]. The usage of active matter RC is in its infancy [62]. Lymburn et al. simulated biologically inspired swarms avoiding an externally controlled predator as a reservoir computer [63], which motivated our work. Note that the converse idea of using a reservoir computer to control a swarm has also been investigated [64], and there is a growing interest in controlling active particles using artificial intelligence [65, 66]. Wang et al. presented an experimental realization of RC with an array of active gold nano-particles driven by a laser [67]. Strong et al. designed RC around active hydrogen ions residing in an electroactive polymer gel, which can be excited by applying an electric field [68]. Two more experimental setups have been conceptually proposed by Jeggle and Wittkowski that leverage driving the active matter agents via spatiotemporally modulated fields [62]: micro-particles in a medium driven by ultrasonic pressure waves, and micro-particles with a symmetry-broken refractive index profile driven by light fields.
The focus of our contribution lies on the characteristics of computing in terms of dynamics, and to encourage further scientific inquiry on the subject. As we will see, computation involves continuously probing matter and utilizing the information gathered by observing responses at various spatial locations, scales, and times. Understanding the nature of these responses is paramount to the essence of computation.
@@
Computation occurs out of thermodynamic equilibrium [69, 70, 71]. Possibilities for retrieving and processing information are constrained by physical laws relating, e.g., dissipation and fluctuations in the system [72]. When (driven) out of thermodynamic equilibrium, complex systems composed of molecules, colloidal, or active particles exhibit collective and correlative features not found in counterparts near equilibrium [73, 74, 75, 76, 77]; the fluctuations internal to these systems are generally broader, and nonlinear coupling mechanisms play a role [78], rendering heterogeneous dynamics. (Adaptive and intelligent systems in nature or elsewhere share key physical markers of nonequilibrium [79, 80, 81, 82].) Dynamics enacting at different scales of time and space is thus a hallmark feature, which, we propose, should indicate the effectiveness of a computing system.
@@
In this paper, we use an active matter model externally driven far from equilibrium by a chaotically varying signal. This model serves as the reservoir computing substrate to predict the future state of the external signal. We employ particle-based simulations, coarse-grained using Gaussian observation kernels, and a readout layer trained using Ridge regression (Sect. II), following the design of Lymburn et al. [63]. We conduct extensive physical parameter scans (Sect. III).
@@
These scans reveal a novel nonequilibrium dynamical regime that demonstrates robustly optimal performance (Sect. III.1), situated near a critical damping point in the intrinsic relaxation dynamics controlled by microscopic forces with the environment. Contrary to previous beliefs, optimal computing is not observed near a gas-droplet phase transition regime of a dynamically rich “critical” or “boiling” swarm, as suggested by Lymburn et al. [63]. Instead, our discovered regime achieves approximately 20 % better predictive performance than previously reported best values.
@@
As we find, the success of computing capabilities is powerfully associated with the intrinsic damping behavior of the system, which is a reflection of the microscopic dynamics (Sect. III.2). Under driving, a complex interplay of coherent and incoherent types of dynamics is revealed in dynamical autocorrelation functions; a system needs to be highly adaptive and dynamically susceptible to the chaotically changing environment. Abrupt, recurring effective mechanisms at mesoscopic scales appear to emerge when information processing becomes both precise and well-generalizing to a task (Sect. III.3). The essence of optimal computation is captured by the intensity and reach of velocity fluctuations (Sect. III.4), serving as physical proxies.
@@
We vary prediction tasks for our reservoir computer and experiment with different dynamic classes of chaotic driving (Sect. III.5). Moreover, we test RC with only one- or two-particle substrates (Sect. III.6), eliminating collective effects on the level of particles, i.e., the theoretical minimum of active matter RC in this model.
@@
The roots of optimal computation thus embody generic features in the mechanisms involved, appear non-specific to the particular input signal, and hint at a generalizing robustness of physical systems that are performant in many scenarios; we synthesize results at the end of the paper (Sect. IV).
@@ II.2 Injecting information of a chaotic time-series into the system via external driving
So far, we described an active matter model that does not receive any external information. To perform reservoir computing and inject time series information into the swarm, we introduce an additional particle called driver (red spiked ball in Fig. 1). It is called the driver because it drives the swarm out of its steady state. We then model a repulsive force between this driver particle and the swarm. The repulsive force
@@
is induced by the driver and experienced by an agent $i$. A cut-off radius $r_{d}$ is in place modeled by the Heaviside step function ($\theta(r_{d}-\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|)=1$ if $\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|\leq r_{d}$, and 00 otherwise). This term is added to the total force sum in Eq. 5 with a term $K_{d}\mathbf{F}_{d,i}$ that enters the $\tanh$ function in Eq. 6. The driver particle itself does not experience a force by the agents; its trajectory is fully determined by an external input time series. This time series is to be processed by the swarm reservoir computer to predict its future state.
Chaotic attractors have been shown to serve as useful benchmarking problems for prediction [83, 31, 84, 85]. They are challenging in prediction tasks because small changes in the initial conditions lead to exponentially different states due to their deterministically chaotic nature. We choose the famous chaotic Lorenz-63 attractor [86] as the main benchmark time series. In addition, chaotic attractors from different dynamical classes (as classified in Ref. [31]) are chosen to assess the performance of our active matter reservoir computer in different scenarios: the attractors Hénon–Heiles [87] (twists), Rössler [88] (rotations), Chua [89] (lobes), and Lorenz96 [90] (ripples). We report on the resulting phenomenological differences of active matter RC for these different attractors in Sect. III.5.
@@ II.4 Readout layer training and prediction
To make a prediction using the reservoir computing approach we train a linear readout layer $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times 3M}$. It connects via weights the $3M$ observations of the swarm at time $t$ with future states of the two-dimensional input time series (the driver trajectory $\mathbf{x}_{d}$) at a time $t+\Delta t_{\mathrm{pred}}$. To train the readout layer we write the input time series
@@
as sequence of driver positions. Then we remove the initial and final $\Delta t_{\mathrm{pred}}$ time steps from $\mathbf{Y}_{\mathrm{target}}$ and $\mathbf{X}^{\mathrm{train}}$, respectively, to align the observations at time $t$ with future states of the time series at $t+\Delta t_{\mathrm{pred}}$. We compute $\mathbf{W}_{\mathrm{out}}$ in the overdetermined linear system
@@ II.5 Observables for driven nonequilibrium soft matter systems
averaged over all agents $N$ and all $N_{\mathcal{T}}$ time windows $\mathcal{T}_{t_{0,j}}=[t_{0,j},t_{0,j}+\tau_{\text{max}})$ within a total number of simulation steps $T$, for a given time lag $\tau\in\mathcal{T}$. Windows $\mathcal{T}$ do not overlap and are separated by a random waiting time drawn uniformly from $\mathcal{T}$ each time.
@@ II.6 Simulation and reservoir computing set-up
We vary the speed-controller parameters $K_{sc},s\in[10^{-5};10^{2}]$ in Sect. III.1, III.4, III.5 and III.6; the alignment force parameters $K_{a}\in[10^{-3};10^{1}]$ and $r_{a}\in[10^{-1};10^{1}]$ for two speed-controller settings $(K_{sc},s)=(2.0,10.0)$ and $(0.02069,0.0483)$ in Sect. III.7; and the homing force strength $K_{h}\in[10^{-3};10^{3}]$ and the speed-controller strength $K_{sc}\in[10^{-5};10^{2}]$ for a fixed target agent speed $s=0.04833$ in Sect. III.8. These parameter scans are performed with 20 values spaced logarithmically for each parameter and plotted in contour plots using linear interpolation (Gouraud shading). All figures show data using testing run initial conditions if not mentioned otherwise; figures showing undriven steady-state systems show data using training run initial conditions. All other parameters are fixed throughout this work unless stated otherwise.
@@
To perform reservoir computing, we simulate our active matter system in two runs: training and testing. For each run we choose different initial conditions for the chaotic input time series, leading to exponentially different trajectories due to their deterministically chaotic nature. Each run consists of an initial burn-in (equilibration) time of 1,000 steps that are neither recorded nor considered for the readout layer training and 50,000 recorded steps for the main simulation. After training the readout layer using the Gaussian kernel observations as described in Sect. II.3 and Sect. II.4, we predict $\Delta t_{\text{pred}}=25$ fixed integration time steps ahead by default at each time point in a non-autoregressive fashion, which corresponds to about half a Lyapunov time for the Lorenz attractor. The predicted time ahead for the other attractors varies and is adjusted to predict Lorenz-63-equivalent Lyapunov times ahead.

### v2

@@ Abstract
Information processing abilities of active matter are studied in the reservoir computing (RC) paradigm to infer the future state of a chaotic signal. We uncover an exceptional regime of agent dynamics that has been overlooked previously. It appears robustly optimal for performance under many conditions, thus providing valuable insights into computation with physical systems more generally. The key to forming effective mechanisms for information processing appears in the system’s intrinsic relaxation abilities. These are probed without actually enforcing a specific inference goal. The dynamical regime that achieves optimal computation is located just below a critical damping threshold, involving a relaxation with multiple stages, and is readable at the single-particle level. At the many-body level, it yields substrates robustly optimal for RC across varying physical parameters and inference tasks. A system in this regime exhibits a strong diversity of dynamic mechanisms under highly fluctuating driving forces. Correlations of agent dynamics can express a tight relationship between the responding system and the fluctuating forces driving it. As this model is interpretable in physical terms, it facilitates re-framing inquiries regarding learning and unconventional computing with a fresh rationale for many-body physics out of equilibrium.
@@ I Introduction
The reservoir computing (RC) paradigm is ideal for exploring these basic questions, as it exploits directly dynamical systems for information processing and inference (in real-time) [25, 26, 4]. The power of RC for scientific inquiry into physical computation is its generic and straightforward setup: Information of a low-dimensional input (e.g. a time series) is injected into the reservoir computing substrate, which is a high-dimensional nonlinear dynamical system that enters different dynamical states depending on the input signal and its history [27]. In the RC paradigm, the reservoir substrate is optimized based on heuristic global properties (such as the spectral radius in Echo State Networks). In the canonical case of utilizing a digital neural network, substrates are simply initialized in a random way. The readout layer [28] typically consists of a linear model trained using Ridge regression. This separation between training the nonlinear substrate and a simple readout layer constitutes the original discovery of RC as a method [25, 26, 4], avoiding costly backpropagation through time [28].
RC renders state-of-the-art performance for time-series forecasting [29, 30, 31, 32], in particular, chaotic time series prediction [30]. Despite the plethora of other machine learning techniques available for time-series forecasting (nonlinear vector autoregression (“next-generation RC”) [33], NBEATS [34], NHiTS [35], transformer models [36] or LSTMs [37]), the highly generic method of RC using e.g. Echo State Networks has often remained an architecture of choice. It requires far less training data [31, 32] and allows for predictions several characteristic times of the chaotic system (Lyapunov times) ahead, in a closed-loop autoregressive fashion [31]. It is considered “model-free” in the sense that it can estimate properties of the input without the need for an explicit model thereof [38]. It can be implemented using generic substrates, e.g., enabling highly promising photonics-based machine learning [39, 40, 41, 42, 43].
@@
However, many questions remain unresolved: For example, what is the optimal mix of nonlinearity and linearity in responses? [48] Do memory capacity and nonlinearity present trade-offs? [49, 48] What is the role of the time delay in response? [50, 51, 41, 40] How can input information be effectively injected into [33] and read out from [52] the system for more efficient RC? And how can interpretable frameworks be developed to integrate theory with practice and enhance the prevalence of RC? [53, 33].
@@
Yet, for reservoir computation in materio, deeper physical concepts remain hidden beneath the surface. For example, how do nonequilibrium phase diagrams and dissipative as well as correlative behavior relate to computational properties? Understanding this fundamental connection would be valuable for recognizing the true feasibility of general neuromorphic computing systems [19, 54, 12], and connect different realms of matter-based and digital computing, both conceptually and practically [55]. Artificial neural networks are usually employed as the model of RC, to develop mathematical theory; they are not particularly interpretable in these terms, essentially functioning as black-boxes [56]. Thus, a different paradigm for thinking about computation and learning is required when dealing with physical substrates [1, 13].
@@
RC has been demonstrated for soft matter, though extremely rarely, and developments remain in an early stage: Reported were cases of gel materials, soft macroscopic bodies, and biological cell cultures [57, 58, 59]. In one exotic example, water waves in a bucket were used as a proof-of-principle [60]. The usage of active matter RC is in its infancy [61]. Lymburn et al. simulated biologically inspired swarms avoiding an externally controlled predator as a reservoir computer [62], which motivated our work. Note that the converse idea of using a reservoir computer to control a swarm has also been investigated [63], and there is a growing interest in controlling active particles using artificial intelligence [64, 65, 66]. Wang et al. presented an experimental realization of RC with an array of active gold nano-particles driven by a laser [67]. Strong et al. designed RC around active hydrogen ions residing in an electroactive polymer gel, which can be excited by applying an electric field [68]. Two more experimental setups have been conceptually proposed by Jeggle and Wittkowski that leverage driving the active matter agents via spatiotemporally modulated fields [61]: micro-particles in a medium driven by ultrasonic pressure waves, and micro-particles with a symmetry-broken refractive index profile driven by light fields.
@@
Physical reservoir computing involves continuously probing matter and using the information gathered by observing responses at various spatial locations, scales, and times. The nature of these responses is the essence of computation, which occurs out of thermodynamic equilibrium [69, 70, 71]. When (driven) out of thermodynamic equilibrium, complex systems composed of molecules, colloidal, or active particles exhibit collective and correlative features not found in counterparts near equilibrium [72, 73, 74, 75, 76, 77, 78, 79]. The fluctuations internal to these systems are generally broader, and nonlinear coupling mechanisms play a role [80], rendering heterogeneous dynamics. These can stimulate the formation of new structures [76]. Notably, intelligent systems in nature or elsewhere share these and related key physical markers [81, 82, 83, 84].
@@
In this paper, we use an active matter model externally driven far from equilibrium by a chaotically varying signal. This active system serves as the reservoir computing substrate to predict the future state of the external signal. We employ particle-based simulations. The physical information is coarse-grained using Gaussian kernels in an observation layer, and a readout layer is trained using Ridge regression (Sect. II), following the design of Lymburn et al. [62]. We conduct extensive physical parameter scans (Sect. III).
@@
These scans reveal a novel nonequilibrium dynamical regime that robustly establishes optimal performance in the RC (Sect. III.1). It is situated near a critical damping point in the intrinsic relaxation dynamics. These individual-particle, non-collective properties give rise to abilities essential for reliable computation. Favorable collective qualities emerge from them. Contrary to previous beliefs, optimal computing is not observed near a gas-droplet phase transition regime of a dynamically rich “critical” or “boiling” swarm, as suggested by Lymburn et al. [62]. Instead, our discovered regime achieves approximately 20 % better predictive performance (using the same metric) than previously reported best values.
@@
As we find, the success of computing capabilities is powerfully associated with the intrinsic relaxation behavior of the system, which is a direct effect of the microscopic dynamics (Sect. III.2). Under driving, a complex interplay is revealed in dynamical autocorrelation functions, including an adaptation to the chaotically changing environment. Intriguing effective mechanisms at mesoscopic scales appear (Sect. III.3). The intensity and reach of velocity fluctuations (Sect. III.4) serve as key physical proxies for computational quality.
@@
We vary prediction tasks for our reservoir computer and experiment with different dynamic classes of chaotic driving (Sect. III.5). Moreover, we test RC with only one- or two-particle substrates (Sect. III.6), eliminating collective effects, i.e., the theoretical minimum of active matter RC in this model.
@@
The roots of optimal computation thus embody generic features in the microscopic dynamical mechanisms involved. These are non-specific to the particular chaotic input signal and suggest a deeper reason for the characteristic robustness of certain physical substrates, which perform well across various scenarios. We synthesize results at the end of the paper (Sect. IV).
@@ II.2 Injecting information of a chaotic time-series into the system via external driving
So far, we have described an active matter model that does not receive any external information. To perform reservoir computing and inject time series information into the swarm, we introduce an additional particle called driver (red spiked ball in Fig. 1). It is called the driver because it drives the swarm out of its steady state. We then model a repulsive force between this driver particle and the swarm. The repulsive force
@@
is induced by the driver and experienced by an agent $i$. A cut-off radius $r_{d}$ is in place modeled by the Heaviside step function ($\theta(r_{d}-\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|)=1$ if $\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|\leq r_{d}$, and 00 otherwise). This term is added to the total force sum in Eq. 5 with a term $K_{d}\mathbf{F}_{d,i}$ that enters the $\tanh$ function in Eq. 6. The driver particle itself does not experience a force from the agents; its trajectory is fully determined by an external input time series. This time series is to be processed by the swarm reservoir computer to predict its future state.
Chaotic attractors have been shown to serve as useful benchmarking problems for prediction [85, 31, 86, 87]. They are challenging in prediction tasks because small changes in the initial conditions lead to exponentially different states due to their deterministically chaotic nature. We choose the famous chaotic Lorenz-63 attractor [88] as the main benchmark time series. In addition, chaotic attractors from different dynamical classes (as classified in Ref. [31]) are chosen to assess the performance of our active matter reservoir computer in different scenarios: the attractors Hénon–Heiles [89] (twists), Rössler [90] (rotations), Chua [91] (lobes), and Lorenz-96 [92] (ripples). We report on the resulting phenomenological differences of active matter RC for these different attractors in Sect. III.5.
@@ II.4 Readout layer training and prediction
To make a prediction using the reservoir computing approach, we train a linear readout layer $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times 3M}$. It connects via weights the $3M$ observations of the swarm at time $t$ with future states of the two-dimensional input time series (the driver trajectory $\mathbf{x}_{d}$) at a time $t+\Delta t_{\mathrm{pred}}$. To train the readout layer, we write the input time series
@@
as a sequence of driver positions. Then we remove the initial and final $\Delta t_{\mathrm{pred}}$ time steps from $\mathbf{Y}_{\mathrm{target}}$ and $\mathbf{X}^{\mathrm{train}}$, respectively, to align the observations at time $t$ with future states of the time series at $t+\Delta t_{\mathrm{pred}}$. We compute $\mathbf{W}_{\mathrm{out}}$ in the overdetermined linear system
@@ II.5 Observables for driven nonequilibrium soft matter systems
averaged over all agents $N$ and all $N_{\mathcal{T}}$ time windows $\mathcal{T}_{t_{0,j}}=[t_{0,j},t_{0,j}+\tau_{\text{max}})$ within a total number of simulation steps $T$, for a given time lag $\tau\in\mathcal{T}$. Windows $\mathcal{T}_{t_{0,j}}$ do not overlap and are separated by a random waiting time drawn uniformly from $\mathcal{T}$ each time.
@@ II.6 Simulation and reservoir computing set-up
We vary the speed-controller parameters $K_{sc},s\in[10^{-5};10^{2}]$ in Sect. III.1, III.4, III.5 and III.6; the alignment force parameters $K_{a}\in[10^{-3};10^{1}]$ and $r_{a}\in[10^{-1};10^{1}]$ for two speed-controller settings $(K_{sc},s)=(2.0,10.0)$ and $(0.02069,0.0483)$ in Sect. III.7; the homing force strength $K_{h}\in[10^{-3};10^{3}]$ and the speed-controller strength $K_{sc}\in[10^{-5};10^{2}]$ for a fixed target agent speed $s=0.04833$ in Sect. III.8. These parameter scans are performed with 20 values spaced logarithmically for each parameter and plotted in contour plots using linear interpolation (Gouraud shading). All figures show data using testing run initial conditions if not mentioned otherwise; figures showing undriven steady-state systems show data using training run initial conditions. All other parameters are fixed throughout this work unless stated otherwise.
@@
To perform reservoir computing, we simulate our active matter system in two runs: training and testing. For each run we choose different initial conditions for the chaotic input time series, leading to exponentially different trajectories due to their deterministically chaotic nature. Each run consists of an initial burn-in (equilibration) time of 1,000 steps that are neither recorded nor considered for the readout layer training and 50,000 recorded steps for the main simulation. After training the readout layer using the Gaussian kernel observations as described in Sect. II.3 and Sect. II.4, we predict $\Delta t_{\text{pred}}=25$ fixed integration time steps ahead by default at each time point in a non-autoregressive fashion, which corresponds to $\approx 0.45283$ Lyapunov times for the Lorenz attractor. The predicted Lyapunov time ahead for the other chaotic attractors is equivalent to that of the Lorenz-63 system.